# Mitigate Position Bias in Large Language Models via Scaling a Single Dimension

### v1

@@ 4.2 Main Results
Table 1: Performance of different methods with different models on NaturalQuestions (20 docs) [10] and KV retrieval (140 KV pairs) [10] dataset.
@@
Table 2: Performance of different methods with different models on LongBench [31].
@@
LLaMA-2-7b-chat 28.9 29.7 6.6 26.3 10.2 12.2 19.0
LLaMA-2-7b-chat w/ Ms-PoE 29.8 31.7 10.5 26.7 6.4 13.2 19.7
LLaMA-2-7b-chat w/ Ours 29.2 29.3 9.7 25.0 18.9 20.8 22.1
LLaMA-2-13b-chat 21.4 14.6 11.2 26.1 4.7 16.9 15.8
LLaMA-2-13b-chat w/ Ms-PoE 20.8 15.4 12.7 27.3 3.1 15.7 15.8
LLaMA-2-13b-chat w/ Ours 30.6 9.6 10.8 25.7 27.6 18.7 20.5
Vicuna-7b-v1.5-16k 30.2 21.6 7.2 26.7 9.4 21.2 19.4
Vicuna-7b-v1.5-16k w/ Ms-PoE 32.3 24.2 8.3 28.0 9.8 22.2 20.8
Vicuna-7b-v1.5-16k w/ Ours 27.1 22.1 11.2 26.1 16.7 20.2 20.6
Vicuna-13b-v1.5-16k 31.1 33.8 21.2 26.2 21.6 23.8 26.3
Vicuna-13b-v1.5-16k w/ Ms-PoE 34.5 33.1 16.0 27.5 21.0 25.0 26.2
Vicuna-13b-v1.5-16k w/ Ours 30.1 35.1 25.0 25.8 27.0 24.7 27.9
Mistral-7b-Instruct-v0.2 37.8 28.5 49.7 28.8 49.9 44.0 39.8
Mistral-7b-Instruct-v0.2 w/ Ms-PoE 41.7 22.2 38.4 24.9 14.0 19.5 26.8
Mistral-7b-Instruct-v0.2 w/ Ours 38.4 30.4 49.8 29.4 51.4 45.3 40.8
@@
Qwen1.5-7b-chat 46.4 39.5 38.4 22.3 39.9 44.6 38.5
Qwen1.5-7b-chat w/ Ms-PoE 42.0 41.5 30.3 25.7 43.2 41.4 37.4
Qwen1.5-7b-chat w/ Ours 45.8 38.8 38.5 22.1 40.0 48.1 38.9
MPT-30b-chat 27.9 21.9 7.5 25.7 18.8 16.7 19.7
@@
MPT-30b-chat w/ Ours 29.4 19.5 6.7 25.8 23.0 21.2 20.9
Tables 1, 2, and 6 present the performance of various methods in different benchmarks. Several observations and conclusions can be drawn: 1) Our method consistently improves overall performance at different positions, with increases of up to 9.3%, 15.2%, and 4.7% in NQ, KV retrieval, and LongBench, respectively, except for LLaMA-2-13B in KV retrieval. Additionally, compared to the SoTA method Ms-PoE, our method shows significant improvements of up to 6.3%, 97.5%, and 14% in NQ, KV retrieval, and LongBench. The poor performance of Ms-PoE in KV retrieval can be attributed to the interpolation causing information loss. 2) Our method effectively enhances LLMs’ understanding of information located in the middle and latter parts of the prompt. For key information at the beginning of the prompt, performance is comparable to baselines. Considering only the average performance of the last four positions, our method’s improvements over the original increase to 11.3% and 16.8% in NQ and KV retrieval, respectively, and over Ms-PoE increase to 8.7% and 97.5% in NQ and KV retrieval, respectively. 3) Our approach is effective not only for RoPE models but also for context window extended models like Vicuna-16K, which already readjust RoPE [34]. Additionally, our method can be adapted to different position embeddings, such as Alibi [22] models like MPT, resulting in improvements of 2.6%, 15.2%, and 1.2% in NQ, KV retrieval, and LongBench, respectively. 4) Our method demonstrated varying degrees of improvement across different tasks, with the most significant increases being 22.9% in few-shot learning tasks, 8.6% in code tasks, 4% in synthetic tasks, 9.2% in single document QA tasks, and 1.9% in multi-document QA tasks. In summarization tasks, performance was nearly on par with the original results. 5) Our method does not disrupt the necessary position information in LLMs, as detailed in Appendix H.
@@ 4.3 Analysis
Figure 5: Attention distribution and performance when scaling dimension 2393 of Vicuna-7b-v1.5-16k with different scale factors on KV retrieval [10] of 100 KV pairs.
The scaling factor directly controls the degree and direction of the impact of position hidden states on position bias. As shown in Figure 5, when the scaling factor is positive, the model exhibits a clear bias towards focusing more on the beginning, while when negative, this bias shifts to focusing more on the end. The factor between 0.5 and -1 leads attention to the most balanced distribution, meanwhile, the improvement in accuracy also reaches its peak. This result proves that the positional hidden states we scaled can indeed influence the bias of LLMs towards focusing excessively on the beginning. By adjusting the coefficients appropriately, this bias can be effectively mitigated.
@@
Table 3: Average performance of different ground-truth positions using different methods on NaturalQuestions multi-document QA dataset (20 docs) [10].
@@
To evaluate the contributions of different components in our method, we introduce the following sets for the ablation study: (1) Ours w/o monotonicity, w/o smoothness, and w/o validation set, which adjust the search algorithm by not considering these three indicators, respectively (details in Appendix C.2). (2) Ours w/ scale 2 dimensions, which modifies the top-2 positional hidden states simultaneously. (3) Ours w/ modify last 16 tokens and w/ modify all tokens, which adjust the range of tokens affected by the scaling operation in Equ.(5).
@@ 5 Related Works
Recent research has focused on expanding the context window size of LLMs. The main approaches include: 1) Staged pre-training [35, 36]: Gradually increasing the context window size during training. 2) Modifying or interpolating position embeddings [22, 34, 37, 38]. 3) Utilizing external memory modules for context storage [39, 40]. 4) Expanding computations across multiple devices in a distributed manner [41]. While these methods address context window expansion, their impact on positional bias in downstream tasks has yet to be discussed.
@@
Although LLMs incorporate explicit positional information through methods like RoPE [21] or Alibi [22], studies such as [10, 16] have found that LLMs exhibit varying degrees of position bias, referred to as "lost in the middle." Recent works aimed to mitigate this issue and improve LLM performance in long-context scenarios can be categorized as follows: 1) RoPE-based methods: These approaches modify the RoPE computation process to alleviate long-distance information decay, including Attention Bucket [20], which uses an ensemble of multiple RoPE bases to mitigate position bias, and Ms-PoE [15], which dynamically interpolates with a small coefficient for different heads. 2) SFT-based methods [14, 18, 19]: These methods construct data with more diverse key information distributions or employ system2think SFT tasks to mitigate position bias. They require further training of the model. 3) Attention mask-based methods [42]: These methods modify attention mechanisms, including Attention Transition [43], which redirects attention to significant parts of the context and Stable Mask [44], which introduces pseudo attention into the causal mask, ensuring stable attention distribution when facing lengthy texts. 4) Prompt-based methods [45, 46]: These methods introduce an external module to reorder or compress information in the prompt, thereby mitigating position bias.
@@ 6 Conclusion
Appendix A Limitations
For different LLMs, an additional search cost is required to determine the positional hidden states including indentify top-K positional hidden and evaluate on the small size validation dataset, which takes approximately 10 minutes for 7B level LLMs on a single A100 GPU.
Due to the difficulty in fully explaining the impact of some dimension to LLM’s behavior, the top-1 dimension returned by the searching algorithm may have the optimal performance on various tasks. Sometimes the optimal dimension is in the top-3 results.
Because of the recalculation of the query and key states, the additional time cost will increase linearly with the length of the input sequence.
Appendix B Broader Impacts
Our methods explore the intrinsic causes of position bias in LLMs and propose a way to alleviate this bias by pruning positional hidden states. This improves the inference capabilities of LLMs, making them more applicable to a wider range of scenarios, especially long-context LLMs and more complex applications. Additionally, our work promotes further research on the deep relationships between the causal mask, hidden states, attention weights and position bias in LLMs, aiding in the iterative development of related technologies.
Appendix C Experiment Details
C.1 Datasets Details
We choose NaturalQuestion Multi-document QA and Key-Value Retrieval datasets used in "lost in the middle" paper [10] to evaluate the degree to which our method alleviates position bias. NaturalQuestion Multi-document QA require the model to answer the question based on one key information document which is inserted in a long context consisting of many irrelevant documents. And Key-Value Retrieval needs the model to retrieve the value corresponding to the given key from a list consisting of hundreds of Key-Value pairs. These two datasets are both classic in-context tasks which aim to evaluate the differences of model performance when key information is located at different positions in the context. The evaluation metric is accuracy, based on whether the model’s response contains a string of the correct answer. In addition, we evaluate our method’s improvements across multi task types, using LongBench [31], a benchmark for bilingual, multitask, and comprehensive assessment of long context understanding capabilities of LLMs. It contains six major categories, covering single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks and code completion. The evaluation metrics are: F1 for single-document QA and multi-document QA, Rouge-L for summarization, accuracy (exact match) for few-shot learning and synthetic tasks, and edit similarity for code completion. During inference, since the original context may sometimes be too long, the input sequences will be truncated in the middle part to avoid exceeding the context window of the model.
C.2 Additional Implemention Details
@@
When applying Ms-PoE [15] to mistral-7b [27] with its default parameters (minimal scale factor is 1.2 and maximal is 1.8), we found the model fail to generate normal responses, so we set the maximal scale factor to 1.2, under which Ms-PoE [15] is equal to PI [34] with scale factor 1.2.
@@
C.3 Scaled Dimensions Details
@@ C.4 Inference Latency
Table 5 shows the running time of LLaMA-2-7b-chat with different methods in the KV retrieval dataset consisting of 500 samples with average length of about 10,000, and the multi-document QA dataset consisting of 500 samples with average length of about 3,300. Our method requires recompute the query and key states, thus inevitably requires more time compared to baseline, but the cost is within an acceptable range. In contrast, Ms-PoE [15] need to compute the attention weights twice, resulting in a doubling of time consumption.
Appendix D Obtain Attention to Key Information
@@
Appendix E Attention v.s. Performance
(a) Attention
(b) Attention w/ modification
(c) Accuracy
(d) Accuracy w/ modification
As illustrated in Figure 6, when we manually multiply all the attention weights to the tokens belong to the key information (here we only choose the 25th KV pair, as shown in Figure 6b) by 2 in the model’s forward process on the KV retrieval task, the corresponding retrieval accuracy of the 25th KV also show improvements, while the other parts mainly keep unchanged, as shown in Figure 6d. This result proves that the attention weights for the key information is positively correlated with the retrieval accuracy.
Appendix F How We Modify Causal Mask and Position Embedding in KV Retrieval
@@
Appendix G Perturbation on Causal Mask and Position Embedding
@@
Figure 9: We performed perturbation experiments on the causal mask and position embedding (PE), showing the dimension 213 of hidden states of Mistral-7b [27] using randomly synthesized corpus as input.
@@
Appendix H Does this Method Compromise the Ability to Perceive Positional Information?
Table 6: Performance of difference models on position-sensitive task-timeline reorder [11].
Model Accuracy
Vicuna-7b-v1.5-16k 20.83
Vicuna-7b-v1.5-16k w/ Ours 20.83
Qwen1.5-7b-chat 28.13
Qwen1.5-7b-chat w/ Ours 28.13
Mistral-7B-Instruct-v0.2 18.75
Mistral-7B-Instruct-v0.2 w/ Ours 19.79
To prove our method is harmless for position-sensitive tasks though eliminating some positional information, we conduct the timeline reorder task from LooGLE [11], whose objective is to arrange the events in accordance with their chronological sequence as dispersed throughout the extensive text. The results in Table 6 proved our methods will not impair model’s performance on position-sensitive tasks. This also reflects that the positional information we eliminate may not be necessary for the model to function well.
Appendix I Attention Distribution Layer-wise and Head-wise
@@
Figure 11 shows the head-wise situation of layer 15. We can see actually only a portion of attention heads exhibit diagonal patterns, which may correspond to retrieval heads [47]. The attention distribution in these heads also shows a pattern corresponding "loss in the middle", being larger at the beginning or end while significantly smaller at the middle.
@@
Appendix J Positional Hidden States Visualization

### v2

@@ 4.2 Main Results
Table 1: Performance of different methods with different models on NaturalQuestions (20 docs) (Liu et al., 2024c) and KV retrieval (140 KV pairs) (Liu et al., 2024c) dataset.
@@
LLaMA-2-7b-chat 28.9 29.7 6.6 26.3 61.2 47.1 33.3
LLaMA-2-7b-chat w/ Ms-PoE 29.8 31.7 10.5 26.7 61.0 48.1 34.6
LLaMA-2-7b-chat w/ Ours 29.2 29.3 9.7 25.0 61.6 46.9 33.6
LLaMA-2-13b-chat 21.4 14.6 11.2 26.1 61.5 39.8 29.1
LLaMA-2-13b-chat w/ Ms-PoE 20.8 15.4 12.7 27.3 62.8 36.3 29.2
LLaMA-2-13b-chat w/ Ours 30.6 9.6 10.8 25.7 62.6 43.2 30.4
Vicuna-7b-v1.5-16k 30.2 21.6 7.2 26.7 53.9 40.5 30.0
Vicuna-7b-v1.5-16k w/ Ms-PoE 32.3 24.2 8.3 28.0 55.2 43.1 31.8
Vicuna-7b-v1.5-16k w/ Ours 27.1 22.1 11.2 26.1 55.0 40.2 30.3
Vicuna-13b-v1.5-16k 31.1 33.8 21.2 26.2 62.0 39.8 35.7
Vicuna-13b-v1.5-16k w/ Ms-PoE 34.5 33.1 16.0 27.5 64.5 37.6 35.5
Vicuna-13b-v1.5-16k w/ Ours 30.1 35.1 25.0 25.8 63.5 41.7 36.9
Mistral-7b-Instruct-v0.2 37.8 28.5 49.7 28.8 65.3 52.9 43.8
Mistral-7b-Instruct-v0.2 w/ Ms-PoE 41.7 22.2 38.4 2.8 23.8 19.5 24.7
Mistral-7b-Instruct-v0.2 w/ Ours 38.4 30.4 49.8 29.4 64.8 52.9 44.3
@@
Qwen1.5-7b-chat 46.4 39.5 38.4 22.3 56.4 50.2 42.2
Qwen1.5-7b-chat w/ Ms-PoE 42.0 41.5 30.3 25.7 46.5 38.0 37.3
Qwen1.5-7b-chat w/ Ours 45.8 38.8 38.5 22.1 57.6 49.6 42.2
MPT-30b-chat 27.9 21.9 7.5 25.7 57.3 39.3 29.9
@@
MPT-30b-chat w/ Ours 29.4 19.5 6.7 25.8 57.6 40.1 29.9
Table 2: Performance of different methods with different models on LongBench (Bai et al., 2024).
Tables 1 and 2 present the performance of various methods in different benchmarks. Several observations and conclusions can be drawn: 1) Our method consistently improves overall performance at different positions, with increases of up to 9.3%, 15.2%, and 4.7% in NQ, KV retrieval, and LongBench, respectively, except for LLaMA-2-13B in KV retrieval. Additionally, compared to the SoTA method Ms-PoE, our method shows significant improvements of up to 6.3%, 97.5%, and 14% in NQ, KV retrieval, and LongBench. The poor performance of Ms-PoE in KV retrieval can be attributed to the interpolation causing information loss. 2) Our method effectively enhances LLMs’ understanding of information located in the middle and latter parts of the prompt. For key information at the beginning of the prompt, performance is comparable to baselines. Considering only the average performance of the last four positions, our method’s improvements over the original increase to 11.3% and 16.8% in NQ and KV retrieval, respectively, and over Ms-PoE increase to 8.7% and 97.5% in NQ and KV retrieval, respectively. 3) Our approach is effective not only for RoPE models but also for context window extended models like Vicuna-16K, which already readjust RoPE (Chen et al., 2023). Additionally, our method can be adapted to different position embeddings, such as Alibi (Press et al., 2022) models like MPT, resulting in improvements of 2.6%, 15.2%, and 1.2% in NQ, KV retrieval, and LongBench, respectively. 4) Our method demonstrated varying degrees of improvement across different tasks, with the most significant increases being 1.5% in few-shot learning tasks, 3.4% in code tasks, 4% in synthetic tasks, 9.2% in single document QA tasks, and 1.9% in multi-document QA tasks. In summarization tasks, performance was nearly on par with the original results. While our method did not significantly improve the average scores overall, it at least demonstrates that it can mitigate position bias without impairing the model’s original capability to handle long context tasks.
@@ 4.3 Analysis
Figure 5: Attention distribution and performance when scaling dimension 2393 of Vicuna-7b-v1.5-16k with different scale factors on KV retrieval (Liu et al., 2024c) of 100 KV pairs.
The scaling factor directly controls the degree and direction of the impact of positional hidden states on position bias. As shown in Figure 5, a positive scaling factor causes the model to focus more on the beginning, while a negative factor shifts the focus towards the end. A factor between 0.5 and -1 leads to the most balanced attention distribution, where accuracy also peaks. These results demonstrate that scaling positional hidden states can influence LLMs’ tendency to focus on the beginning, and by adjusting the coefficients, this bias can be effectively mitigated.
@@
Table 3: Average performance of different ground-truth positions using different methods on NaturalQuestions multi-document QA dataset (20 docs) Liu et al. (2024c).
To evaluate the contributions of different components in our method, we introduce the following sets for the ablation study: (1) Ours w/o monotonicity, w/o smoothness, and w/o validation set, which adjust the search algorithm by not considering these three indicators, respectively (details in Appendix A.2). (2) Ours w/ scale 2 dimensions, which modifies the top-2 positional hidden states simultaneously. (3) Ours w/ modify last 16 tokens and w/ modify all tokens, which adjust the range of tokens affected by the scaling operation in Equ.(5).
@@
Side Effects
We utilized the MMLU dataset (Hendrycks et al., 2020), which assesses general capabilities, and the timeline-reorder dataset (Li et al., 2024a), which is a task sensitive to positional information, to evaluate whether our approach adversely affects the original abilities of the LLM.
Attention v.s. Performance
(a) Attention
(b) Attention w/ modification
(c) Accuracy
(d) Accuracy w/ modification
As shown in Figure 6, when we manually double the attention weights of the key information (in this case, the 25th KV pair, as illustrated in Figure 6b) during the model’s forward pass on the KV retrieval task, the retrieval accuracy for the 25th KV improves, while the accuracy for other parts remains largely unchanged (Figure 6d). This demonstrates that the attention weights for key information are positively correlated with retrieval accuracy.
Does this Method Compromise the Ability to Perceive Positional Information?
Model MMLU Reorder
Vicuna-7b-v1.5-16k 48.22 20.83
Vicuna-7b-v1.5-16k w/ Ours 48.38 20.83
Qwen1.5-7b-chat 60.84 28.13
Qwen1.5-7b-chat w/ Ours 61.43 28.13
Mistral-7B-Instruct-v0.2 60.31 18.75
Mistral-7B-Instruct-v0.2 w/ Ours 60.38 19.79
Table 4: Performance of difference models on MMLU and the timeline reorder task.
To demonstrate that our method does not harm the model’s performance on general or position-sensitive tasks, despite eliminating some positional information, we tested it on two datasets: the MMLU benchmark (Hendrycks et al., 2020) and the timeline reorder task from LooGLE (Li et al., 2024a), which involves arranging events chronologically across extensive text. As shown in Table 4, our method does not impair performance on position-sensitive tasks, indicating that the positional information we remove may not be essential for the model’s functioning.
@@ 5 Related Works
Recent research has focused on expanding the context window size of LLMs through four main approaches: 1) Staged pre-training (Nijkamp et al., 2023; Fu et al., 2024), which gradually increases the context window size during training; 2) Modifying or interpolating position embeddings (Press et al., 2022; Chen et al., 2023; Peng et al., 2024; Ding et al., 2024); 3) Using external memory modules for context storage (Bertsch et al., 2023; Tworkowski et al., 2023); 4) Distributed computation across devices (Liu et al., 2024b). While these methods address context expansion, their impact on positional bias in downstream tasks has not been thoroughly explored.
@@
Despite explicit positional encoding methods like RoPE (Su et al., 2024) and Alibi (Press et al., 2022), LLMs often exhibit position bias, such as the "lost in the middle" phenomenon (Liu et al., 2024c; Kamradt, 2023). Recent efforts to mitigate this bias fall into several categories: 1) RoPE-based methods: These approaches modify the RoPE computation process to alleviate long-distance information decay, including Attention Bucket (Chen et al., 2024), which uses an ensemble of multiple RoPE bases to mitigate position bias, and Ms-PoE (Zhang et al., 2024), which dynamically interpolates with a small coefficient for different heads. 2) SFT-based methods (Junqing et al., 2023; Yu, 2023; An et al., 2024): These methods construct data with more diverse key information distributions or employ system2think SFT tasks to mitigate position bias. They require further training of the model. 3) Attention mask-based methods (He et al., 2024): These methods modify attention mechanisms, including Attention Transition (Gao et al., 2023), which redirects attention to significant parts of the context and Stable Mask (Yin et al., 2024), which introduces pseudo attention into the causal mask, ensuring stable attention distribution when facing lengthy texts. 4) Prompt-based methods (Jiang et al., 2024; Peysakhovich & Lerer, 2023): These methods introduce an external module to reorder or compress information in the prompt, thereby mitigating position bias.
@@ 6 Conclusion
Appendix A Experiment Details
A.1 Datasets Details
We choose NaturalQuestion Multi-document QA and Key-Value Retrieval datasets used in "lost in the middle" paper (Liu et al., 2024c) to evaluate the degree to which our method alleviates position bias. NaturalQuestion Multi-document QA require the model to answer the question based on one key information document which is inserted in a long context consisting of many irrelevant documents. And Key-Value Retrieval needs the model to retrieve the value corresponding to the given key from a list consisting of hundreds of Key-Value pairs. These two datasets are both classic in-context tasks which aim to evaluate the differences of model performance when key information is located at different positions in the context. The evaluation metric is accuracy, based on whether the model’s response contains a string of the correct answer. In addition, we evaluate our method’s improvements across multi task types, using LongBench (Bai et al., 2024), a benchmark for bilingual, multitask, and comprehensive assessment of long context understanding capabilities of LLMs. It contains six major categories, covering single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks and code completion. The evaluation metrics are: F1 for single-document QA and multi-document QA, Rouge-L for summarization, accuracy (exact match) for few-shot learning and synthetic tasks, and edit similarity for code completion. During inference, since the original context may sometimes be too long, the input sequences will be truncated in the middle part to avoid exceeding the context window of the model.
A.2 Additional Implemention Details
@@
When applying Ms-PoE (Zhang et al., 2024) to mistral-7b (Jiang et al., 2023) with its default parameters (minimal scale factor is 1.2 and maximal is 1.8), we found the model fail to generate normal responses, so we set the maximal scale factor to 1.2, under which Ms-PoE (Zhang et al., 2024) is equal to PI (Chen et al., 2023) with scale factor 1.2.
@@
A.3 Scaled Dimensions Details
@@ A.4 Inference Latency
Table 6 shows the running time of LLaMA-2-7b-chat with different methods in the KV retrieval dataset consisting of 500 samples with average length of about 10,000, and the multi-document QA dataset consisting of 500 samples with average length of about 3,300. Our method requires recompute the query and key states, thus inevitably requires more time compared to baseline, but the cost is within an acceptable range. In contrast, Ms-PoE (Zhang et al., 2024) need to compute the attention weights twice, resulting in a doubling of time consumption.
Appendix B Obtain Attention to Key Information
@@
Appendix C How We Modify Causal Mask and Position Embedding in KV Retrieval
@@
Appendix D Perturbation on Causal Mask and Position Embedding
@@
Figure 9: We performed perturbation experiments on the causal mask and position embedding (PE), showing the dimension 213 of hidden states of Mistral-7b (Jiang et al., 2023) using randomly synthesized corpus as input.
@@
Appendix E Attention Distribution Layer-wise and Head-wise
@@
Figure 11 shows the head-wise situation of layer 15. We can see actually only a portion of attention heads exhibit diagonal patterns, which may correspond to retrieval heads (Wu et al., 2024). The attention distribution in these heads also shows a pattern corresponding "loss in the middle", being larger at the beginning or end while significantly smaller at the middle.
@@
Appendix F Positional Hidden States Visualization